Cos(a)+ cos(b)+ cos(c)+ cos(a+b+c)=4cos([a+b]/2)cos([b+c]/2)cos([a+c]/2)

Grade:11

1 Answers

Sanskar sharma 814
6 years ago
Proof=Let the RHS of given equation be EE= 2cos([a+b]/2)cos([a-b]/2)+2cos([a+b+2c]/2)cos([a-b]/2)=2cos([a+b]/2){2cos([a+c]/2)cos([b+c]/2)}=4cos([a+b]/2)cos([b+c]/2)cos([a+c]/2) LHS=RHS(use cos(m) +cos (n) = 2cos([m+n]/2)cos([m-n]/2)